OpenStudy (anonymous):

A hemispherical tank has a total surface area of 243pi ft^2 (this includes the circular base area) Find the volume Please help!!

OpenStudy (anonymous):

The surface area of a sphere is \[ A_{sphere} = 4\pi r^2\] in this case, a hemisphere is just half a sphere. So it is \[A_{hemisphere} =2 \pi r^2 \] So we solve for r. Then we put that into the equation for the volume of a hempshiere, which again is the volumn of half a sphere \[ A_{sphere} = \frac{4}{3}\pi r^3\] \[ A_{hemisphere} = \frac{2}{3}\pi r^3\] Do you need help solving for a variable and plugging it back in?